# Model-agnostic Coreset Selection via LLM-based Concept Bottlenecks

### v1

@@ 4.2 Concept-based score for CS
Next, we describe how to use the concept bottleneck to produce a difficulty score for the samples in the dataset. We start by discussing how we learn the functions $f$ and $g$ described in Sec. 3.2 (see Fig. 1(block 2) for an overview of the method). We use the dot product between the visual embeddings of an image $x$ denoted as $\mathcal{V}_{enc}(x)$ and the bottleneck embedding matrix $E_{C}$ to measure the alignment between the visual features and textual features of the concepts (Yang et al., 2023b; Yan et al., 2023). Concretely, the concept similarity scores for a sample $x$ is computed as
$g(x,E_{C}):=\mathcal{V}_{enc}(x)\cdot E_{C}^{\intercal}.$ (2)
To map the concept similarity scores to predictions in the label space $\mathcal{Y}$, we propose to use a linear (concept bottleneck layer) predictor as $f$. Concretely, the function $f$ with parameters $W\in\mathbb{R}^{N\times N_{C}}$ is given by $f(x;W):=g(x,E_{C})\cdot W^{\intercal}$. We learn the parameters $W$ for the function $f$ using
@@
where $\ell(f(x;W),y)=-\log(f(x;W)_{y})$ is the cross-entropy loss. We define the concept bottleneck layer’s output, $h(x):=f(g(x,E_{C});W^{*}).$ In practice, we learn the optimal $W$ using mini-batch gradient descent by running the optimization for $T$ epochs. To gauge the difficulty of each training sample we use the area under the margin (AUM) (Pleiss et al., 2020) while solving Eq. 3 which quantifies the data difficulty as the margin of a training sample averaged over $T$ training epochs. Concretely, the margin of a sample $(x,y)$ at a training epoch $t$ is $M^{t}(x,y)=h^{t}_{y}(x)-\max_{y^{\prime}\neq y}h^{t}_{y^{\prime}}(x),$ where $h^{t}_{y^{\prime}}(x)$ is the prediction likelihood of the bottleneck layer at epoch $h^{t}$ for class $y^{\prime}$. Thus, AUM (concept-based score) is computed as
@@
Recent works (Pleiss et al., 2020; Zheng et al., 2022, 2024) have demonstrated the effectiveness of AUM for gauging the sample’s difficulty for coreset selection. However, (Zheng et al., 2022, 2024) requires computing AUM for the downstream model by training it on the entire dataset once, which is computationally costly. On the other hand, our method integrates AUM with the training of the linear concept bottleneck layer $h$, which is computationally cheaper (training a linear layer takes only $7$ minutes for Imagenet compared to $8$ hours for training a ResNet-34) than training the downstream model ($\theta$). Moreover, since our method does not use the downstream model in the computation of the difficulty score, the coresets are independent of the model architecture, unlike previous training dynamics-based approaches.
Sampling training examples to form a coreset. After obtaining data difficulty scores, a crucial step is choosing the samples to form the coreset. While many previous works (Toneva et al., 2018; Coleman et al., 2019) have reported encouraging results keeping the most challenging samples (for our concept-based score this means samples with the smallest margin), recent works (Zheng et al., 2022; Sorscher et al., 2022) have shown that this could lead to a catastrophic drop in accuracies after training the downstream model on the coreset, especially when the size of the coreset is small. This is mainly due to poor sample coverage and potentially mislabeled data in the datasets. To remedy this, we use Coverage-centric Coreset Selection (CCS) proposed by (Zheng et al., 2022) (see Alg. 1 in App. C.2) which uses a stratified sampling approach and filters out (potentially) mislabeled samples to form the coreset. This technique has been shown to achieve consistently superior results to the random baselines for various coreset sizes.
@@ 4.3 Concept-based score for label-free CS
Recently, there has also been an interest (Zheng et al., 2024; Maharana et al., 2023; Griffin et al., 2024) in identifying the representative samples from an unlabeled dataset such that 1) we reduce the samples that need to be labeled by humans and 2) we can improve the efficiency of model training by only training the model on a subset of data. Our concept-based score can also be effectively utilized for this task with a simple modification. Similar to previous works (Maharana et al., 2023; Zheng et al., 2024; Sorscher et al., 2022), we assume that we know the number of classes in the datasets. Additionally, we assume that we also know the names of the classes in the datasets. Previous works have demonstrated that VLMs such as CLIP (Radford et al., 2021) achieve excellent zero-shot performance without requiring fine-tuning on specific datasets. We leverage this capability of CLIP models to obtain pseudo-labels for the images in our unlabeled dataset and use them to obtain our difficulty score for each sample
@@
where for an image $x$ in the dataset, $y_{\texttt{pseudo}}=\arg\max_{j\in\mathcal{Y}}\mathcal{V}_{enc}(x)\cdot\mathcal{W}^{\intercal}_{\texttt{zeroshot}}$ where $\mathcal{W}_{\texttt{zeroshot}}\in\mathbb{R}^{N\times d}$ is a matrix with columns defined as $\mathcal{T}_{enc}(s_{j})$ and $s_{j}=$ “a photo of a {$j^{th}$ class name}” for each class $j\in\mathcal{Y}$ (Radford et al., 2021; Wortsman et al., 2022). We use these scores along with CCS to produce the coreset. Similar to (Zheng et al., 2024; Maharana et al., 2023), this coreset is then assumed to be annotated by humans and downstream models ($\theta$) are trained on this annotated coreset.
@@ 5 Experiments
Here, we compare the performance of the models trained on coresets generated by our concept-based score to models trained on coresets selected by various SOTA CS methods for the standard and label-free CS problems.
@@
Needs Training Dynamics Entropy $94.44$ $92.11$ $85.67$ $66.52$ $72.26$ $63.26$ $50.49$ $28.96$ $72.34$ $70.76$ $64.04$ $39.04$
@@
Doesn’t Need Training Dynamics Random $94.33$ $93.40$ $90.94$ $79.08$ $74.59$ $71.07$ $65.30$ $44.76$ $72.18$ $70.34$ $66.67$ $52.34$
@@
Table 2: Comparison of the model’s (RN-18 for CIFAR10/100 and RN-34 for Imagenet) test accuracy after training on coresets, found in a label-free manner, shows that our coresets lead to better performance than Random and those found by Prototypicality and is competitive/better than the coresets found by ELFS (which is dependent on the training dynamics of the downstream model).
@@
Needs Training Dynamics ELFS (SwAV) $95.00$ $94.30$ $91.80$ $82.50$ $76.10$ $72.10$ $65.50$ $49.80$ $73.20$ $71.40$ $66.80$ $53.40$
@@
Doesn’t Need Training Dynamics Prototypicality $94.70$ $92.90$ $90.10$ $70.90$ $74.50$ $69.80$ $61.10$ $32.10$ $70.90$ $60.80$ $54.60$ $30.60$
@@
Datasets, models, and training: We focus on CS for classification tasks on three benchmark datasets. Specifically, we use CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), and Imagenet-1K (Deng et al., 2009) datasets consisting of $50000$, $50000$, and $1.28$ million samples spread across $10$, $100$, and $1000$ classes, respectively. For CIFAR-10/CIFAR-100, we train a ResNet-18 model and for Imagenet we train ResNet-18, ResNet-34, and ResNet-50 models on the coresets for all pruning ratios. To accelerate training on Imagenet, we utilize the training code based on FFCV (Leclerc et al., 2023). We run CS for three trials with different random seeds for all experiments and report the average of these runs in our tables. We report the results for different pruning rates where a pruning rate of $90$% refers to removing $90$% of the samples from the original training dataset.
For generating the concept annotation we use a recently proposed open source model LLaVA (Liu et al., 2023b, a). Extracting the concepts for each class using this model takes a mere 3 seconds per prompt. In Sec. 5.5, we present an ablation study using concepts extracted from different models/methods. For computing the concept similarity scores between the visual and concept bottleneck features we used the CLIP (Radford et al., 2021) model following the previous works (Yun et al., 2022; Yang et al., 2023b; Yan et al., 2023) which showed its effectiveness for this task. Specifically, we used the ViT B-32 CLIP model (Radford et al., 2021). For computing the pseudo-labels in Sec. 4.3 we used a ViT L-14 CLIP model trained on the DataComp-1B dataset (Ilharco et al., 2021). The accuracies of the models trained on the entire training set are $95.44$% and $78.74$% for ResNet(RN)-18 on CIFAR-10/100 and $72.4$% for RN-18, $75$% for RN-34, and $78.4$% for RN-50 on Imagenet. Further experimental details are mentioned in App. C.
@@
CS baselines and methods: We compare our method against various baselines and SOTA methods proposed by previous works. 1) Random: Uniformly select samples from the datasets to form the coreset. RandomFFCV denotes the performance of the models trained on random subsets of Imagenet using the training code based on FFCV (Leclerc et al., 2023). 2) Entropy (Coleman et al., 2019): Selects samples based on entropy which is computed as the uncertainty in the model’s prediction on a sample. 3) Forgetting Score (Toneva et al., 2018): Selects samples based on the forgetting score which is computed as the number of times an example is misclassified after being correctly classified earlier during training of the downstream model. A higher forgetting score indicates a more challenging sample. 4) AUM (Pleiss et al., 2020): Selects samples based on their average margin during training of the downstream model i.e., the difference between the target class and the next highest class across the training epochs. Lower AUM indicates a more challenging sample. For the forgetting score, AUM, and our method, we use CCS (Zheng et al., 2022) for sampling to form the coreset whereas for entropy we select the samples with the highest entropy in the coreset as done in previous works (Coleman et al., 2019; Zheng et al., 2022).
@@
For label-free CS, we use 1) Prototypicality (Sorscher et al., 2022): which first performs k-means clustering in the embedding space of SwAV (Caron et al., 2020) model and ranks samples based on their Euclidean distance to the cluster centers. Samples further away from the cluster center are then used to form the coreset. 2) ELFS (Zheng et al., 2024): estimates the pseudo-labels of the unlabeled samples using a deep clustering approach (using the embedding space of SwAV (Caron et al., 2020) and DINO (Caron et al., 2021)) and forms the coreset using the training dynamics of the downstream model trained on the pseudo-labeled data.
Crucially, SOTA methods such as forgetting score, AUM, and ELFS require training the downstream model on the entire dataset (with true/pseudo labels) first, for CS, unlike our method which is independent of the downstream model. While the Random and the Prototypicality don’t require the downstream model for CS, we show in the following sections that our results are significantly better than these.
5.1 Evaluating performance on CS
@@
Table 1 shows the accuracy of models trained on coresets found by various approaches on the test sets of the three datasets for the standard CS problem (where the dataset is labeled). We find that coresets found by our approach lead to significantly better performance, even at higher pruning rates, compared to the random subsets. Moreover, our method which does not use any information about the training dynamics of the downstream models provides competitive performance to coresets found by the SOTA approaches based on forgetting score and AUM, and even outperforms them on Imagenet for smaller pruning rates.
5.2 Evaluating performance on label-free CS
@@
Table 2 shows the accuracy of models trained on coresets when the training set lacks labels (we report the numbers presented by (Zheng et al., 2024) for all the methods and baselines). The results show that the random subsets are a competitive baseline and even it outperforms Prototypicality (Sorscher et al., 2022). Our results also show that our coresets outperform the random subsets for all pruning rates with the improvements being the most significant at higher pruning rates. Compared to ELFS (Zheng et al., 2024), our method provides competitive performance and even surpasses it for lower pruning rates on Imagenet, without using any information about the downstream model’s architecture or its training dynamics. Thus, our method is effective for this problem with a simple modification of utilizing the zero-shot predictions from CLIP as pseudo-labels.
@@
5.3 Evaluating the transferability and efficiency of CS
Here we evaluate the performance of training downstream models with three different architectures to highlight the effectiveness of our approach for model-agnostic CS. Table 3 shows superior performance than random for ResNet-18, ResNet-34, and ResNet-50 models trained on our coresets for standard and label-free CS for various pruning rates.
Next, we compare the efficiency of our approach at finding coresets compared to approaches relying on training the downstream model on the entire dataset to obtain the training dynamics-based score. Using two A-100 GPUs, our approach can find the coreset in approximately $30$ minutes for the Imagenet dataset giving a $15$x speed up over training dynamics-based approaches. We obtain a similar speed up for CIFAR-10/100 where our method finds the coreset in less than $2$ minutes. Moreover, since our method is agnostic to the architecture of the downstream model we do not need to repeat the CS step for different architectures, unlike other methods which need to train the downstream model with new model architecture to find the best coreset for it.
Table 3: Superior performance of downstream models with different architectures trained on our coresets for Imagenet compared to Random for standard (Ours) and label-free (Ours-LF) CS highlight our method’s effectiveness for downstream model-agnostic CS.
Model Arch. Method Pruning Rates
@@ 5.3 Evaluating the transferability and efficiency of CS
RN-18 Random $71.15_{\tiny\pm 0.23}$ $68.48_{\tiny\pm 0.10}$ $63.15_{\tiny\pm 0.19}$ $44.96_{\tiny\pm 0.50}$
Ours-LF $71.21_{\tiny\pm 0.09}$ $68.77_{\tiny\pm 0.13}$ $63.76_{\tiny\pm 0.09}$ $47.50_{\tiny\pm 0.10}$
Ours $70.94_{\tiny\pm 0.19}$ $69.30_{\tiny\pm 0.84}$ $65.16_{\tiny\pm 0.04}$ $49.57_{\tiny\pm 0.16}$
RN-34 Random $73.37_{\tiny\pm 0.08}$ $71.71_{\tiny\pm 0.10}$ $67.85_{\tiny\pm 0.04}$ $51.29_{\tiny\pm 0.20}$
Ours-LF $73.61_{\tiny\pm 0.08}$ $71.99_{\tiny\pm 0.05}$ $68.42_{\tiny\pm 0.21}$ $53.52_{\tiny\pm 0.06}$
Ours $73.39_{\tiny\pm 0.12}$ $72.34_{\tiny\pm 0.13}$ $69.44_{\tiny\pm 0.17}$ $55.92_{\tiny\pm 0.02}$
RN-50 Random $76.06_{\tiny\pm 0.11}$ $74.44_{\tiny\pm 0.04}$ $70.50_{\tiny\pm 0.02}$ $53.56_{\tiny\pm 0.13}$
Ours-LF $76.54_{\tiny\pm 0.08}$ $74.84_{\tiny\pm 0.03}$ $71.10_{\tiny\pm 0.09}$ $55.22_{\tiny\pm 0.92}$
Ours $76.29_{\tiny\pm 0.10}$ $75.12_{\tiny\pm 0.03}$ $72.05_{\tiny\pm 0.09}$ $58.26_{\tiny\pm 0.46}$
5.4 Visualizing samples with their concept-based scores

### v2

@@ 4.2 Concept-based score for CS
We now describe how to use the concept bottleneck to produce a difficulty score for the samples in the dataset. We start by discussing how we learn the functions $f$ and $g$ described in Sec. 3.2 (see Fig. 2 (block 2)). We use the dot product between the visual embeddings of an image $x$ i.e. $\mathcal{V}_{enc}(x)$ and the bottleneck embedding matrix $E_{C}$ to measure the alignment between the visual and textual features yang2023language ; yan2023learning . Concretely, we compute the concept similarity score for a sample $x$ as
$g(x;E_{C}):=\mathcal{V}_{enc}(x)\cdot E_{C}^{\intercal}.$ (2)
To map the concept similarity score to a prediction in the label space $\mathcal{Y}$, we propose to use a linear (concept bottleneck layer) predictor denoted by $f$. Concretely, the function $f$ with parameters $W\in\mathbb{R}^{N\times N_{C}}$ is given by $f(x;W):=g(x;E_{C})\cdot W^{\intercal}$. We learn the parameters $W$ using
@@
where $\ell(f(x;W),y)=-\log(f(x;W)_{y})$ is the cross-entropy loss. The output of the concept bottleneck layer is defined as, $h(x):=f(g(x;E_{C});W^{*}).$ In practice, we learn $W$ using mini-batch gradient descent by running the optimization for $T$ epochs. We then compute the difficulty of each training sample using area under the margin (AUM) pleiss2020identifying while solving Eq. 3, quantifying the data difficulty as the margin of a training sample averaged over $T$ training epochs. Concretely, margin of a sample $(x,y)$ at an epoch $t$ is $M^{t}(x,y)=h^{t}_{y}(x)-\max_{y^{\prime}\neq y}h^{t}_{y^{\prime}}(x),$ where $h^{t}_{y^{\prime}}(x)$ is the prediction likelihood of the bottleneck layer at epoch $t$ for a class $y^{\prime}$. Thus, AUM (concept-based score) is
@@
Recent works pleiss2020identifying ; zheng2022coverage ; zheng2024elfs have demonstrated the effectiveness of AUM for computing a sample’s difficulty for CS. However, zheng2022coverage ; zheng2024elfs compute AUM for a specific downstream model by training it on the entire dataset first, which is computationally costly. On the other hand, we integrate AUM with the training of a linear layer $h$, training which is significantly cheaper than training the downstream model (training $h$ for 100 epochs takes only $7$ minutes on Imagenet compared to $8$ hours for training a ResNet-34 for 100 epochs). Moreover, since our score is independent of the downstream model, our coresets can be used for any downstream model without change, unlike training dynamics-based approaches that require computing their coresets again for different/new downstream models.
Sampling training examples to form a coreset. After obtaining data difficulty scores, a crucial step is choosing the samples to form the coreset. While many previous works toneva2018empirical ; coleman2019selection have reported encouraging results keeping the most challenging samples (for our concept-based score this means samples with the smallest margin), recent works zheng2022coverage ; sorscher2022beyond have shown that this could lead to a catastrophic drop in accuracies after training the downstream model on the coreset, especially when the size of the coreset is small. This is mainly due to poor sample coverage and potentially mislabeled data in the datasets. To remedy this, we use Coverage-centric Coreset Selection (CCS) proposed by zheng2022coverage (see Alg. 1 in App. C.5) which filters out (potentially) mislabeled samples and uses a stratified sampling approach to form the coreset. This technique has been shown to consistently achieve superior results to the random baselines for various coreset sizes.
@@ 4.3 Concept-based score for label-free CS
Recently, there has been an interest zheng2024elfs ; maharana2023d2 ; griffin2024zero in identifying the representative samples from an unlabeled dataset so as to 1) reduce the samples that need to be labeled/annotated by humans and 2) improve the efficiency of model training by only training the model on a subset of data. Our concept-based score can also be effectively utilized for this task with a simple modification. Similar to previous works maharana2023d2 ; zheng2024elfs ; sorscher2022beyond , we assume that we know the number of classes in the datasets. Additionally, we assume that we also know the names of the classes in the datasets. Previous works have demonstrated that VLMs such as CLIP radford2021learning achieve excellent zero-shot performance without requiring fine-tuning on specific datasets. We leverage this capability of CLIP models to obtain pseudo-labels for our unlabeled dataset and use them to obtain our difficulty score as follows
@@
where for an image $x$ in the dataset $y_{\texttt{pseudo}}=\arg\max_{j\in\mathcal{Y}}\mathcal{V}_{enc}(x)\cdot\mathcal{W}^{\intercal}_{\texttt{zeroshot}}$ where $\mathcal{W}_{\texttt{zeroshot}}\in\mathbb{R}^{N\times d}$ is a matrix with columns defined as $\mathcal{T}_{enc}(s_{j})$ and $s_{j}=$ “a photo of a {$j^{th}$ class name}" for each class $j\in\mathcal{Y}$ radford2021learning ; wortsman2022robust . We use these scores along with CCS zheng2022coverage to produce the coreset. Similar to zheng2024elfs ; maharana2023d2 , the coreset is then assumed to be annotated by humans and models are trained on it.
@@ 5 Experiments
Datasets, models, and training: We focus on CS for classification tasks on three benchmark datasets namely, CIFAR-10, CIFAR-100 krizhevsky2009learning , and Imagenet-1K deng2009imagenet datasets consisting of $50000$, $50000$, and $1.28$ million samples spread across $10$, $100$, and $1000$ classes, respectively. For CIFAR-10/CIFAR-100, we train a ResNet(RN)-18 as a downstream model and for Imagenet we train ResNet-18/34/50 he2016deep , MobileNet sandler2018mobilenetv2 , DenseNet huang2017densely , Wide ResNet zagoruyko2016wide , and ViT dosovitskiy2020image as downstream models on the coresets. We use FFCV leclerc2023ffcv to accelerate training on Imagenet. We run CS for three trials with different random seeds and report the average of these runs in our tables for various pruning rates where a pruning rate of $90$% refers to removing $90$% of the samples.
For generating concepts we use a recently proposed open source model LLaVA liu2023llava ; liu2023improvedllava . For computing the concept similarity scores between the visual and concept features we used the CLIP radford2021learning model (with the ViT B-32 as backbone) following the previous works yun2022vision ; yang2023language ; yan2023learning . For computing the pseudo-labels for label-free CS (in Sec. 4.3) we used a ViT L-14 CLIP model trained on the DataComp-1B dataset ilharco_gabriel_2021_5143773 . Further experimental details are mentioned in App. C.
@@
Needs downstream model training/dynamics Entropy $94.44$ $92.11$ $85.67$ $66.52$ $72.26$ $63.26$ $50.49$ $28.96$ $72.34$ $70.76$ $64.04$ $39.04$
@@
Does not need downstream model training/dynamics Random $94.33$ $93.40$ $90.94$ $79.08$ $74.59$ $71.07$ $65.30$ $44.76$ $72.18$ $70.34$ $66.67$ $52.34$
@@
Table 2: Comparison of the model’s test accuracy after training on coresets, found in a label-free manner, shows that our coresets lead to better performance than Random and Prototypicality. It is also competitive/better than the coresets found by ELFS (which uses training dynamics of the downstream model).
@@
Needs downstream model training/dynamics ELFS (SwAV) $95.00$ $94.30$ $91.80$ $82.50$ $76.10$ $72.10$ $65.50$ $49.80$ $73.20$ $71.40$ $66.80$ $53.40$
@@
Does not need downstream model training/dynamics Prototypicality $94.70$ $92.90$ $90.10$ $70.90$ $74.50$ $69.80$ $61.10$ $32.10$ $70.90$ $60.80$ $54.60$ $30.60$
@@
CS baselines and methods: We compare our method against various baselines and SOTA CS methods. 1) Random: Uniformly select samples from the datasets to form the coreset. RandomFFCV denotes the performance of the models trained on random subsets of Imagenet using FFCV leclerc2023ffcv . 2) Entropy coleman2019selection : Selects samples based on entropy computed as the uncertainty of a model’s prediction on a sample. 3) Forgetting toneva2018empirical : Selects samples based on the forgetting score computed as the number of times a sample is misclassified after being correctly classified earlier during training of the downstream model. A higher forgetting score indicates a more challenging sample. 4) AUM pleiss2020identifying : Selects samples based on their average margin during training of the downstream model i.e., the difference between the target class and the next highest class across the training epochs. Lower AUM indicates a more challenging sample. For forgetting, AUM, and our method, we use CCS zheng2022coverage to form the coreset whereas for entropy we select samples with the highest entropy as done in coleman2019selection ; zheng2022coverage .
@@
For label-free CS, we use 1) Prototypicality sorscher2022beyond : which first performs k-means clustering in the embedding space of SwAV caron2020unsupervised model and ranks samples based on their Euclidean distance to the cluster centers. Samples further away from the cluster center are used to form the coreset. 2) ELFS zheng2024elfs : estimates the pseudo-labels of the unlabeled samples using a deep clustering approach (using the embedding space of SwAV caron2020unsupervised and DINO caron2021emerging ) and forms the coreset using training dynamics of the downstream model trained on the pseudo-labeled data. Crucially, SOTA methods such as forgetting, AUM, and ELFS require training the downstream model on the entire dataset first for CS, unlike our method which is independent of the downstream model. While Random and Prototypicality also don’t require the downstream model for CS, our results show that our approach surpasses them.
5.1 Evaluating performance of our concept-based score for CS
@@
Table 1 shows the accuracy of models trained on coresets found by various approaches on three datasets for the standard CS problem (where the dataset is labeled). Our results show that our coresets lead to significantly better performance, even at higher pruning rates, compared to the random subsets. Moreover, our method which does not have the knowledge of the downstream model or its training dynamics still provides competitive performance to coresets found by the SOTA approaches based on forgetting and AUM, and even outperforms them on Imagenet for smaller pruning rates.
@@
Table 2 shows the accuracy of models trained on coresets when the training set is unlabeled (we report the numbers presented by zheng2024elfs for previous methods). The results show that the random subsets are a competitive baseline for this problem outperforming Prototypicality sorscher2022beyond . Our results also show that our coresets outperform the random subsets for all pruning rates with significant improvements at higher pruning rates. Compared to ELFS zheng2024elfs , our method provides competitive performance and even surpasses it for lower pruning rates on Imagenet, without using any information about the downstream model’s architecture or its training dynamics highlighting its effectiveness.
Next, evaluate the transferability of the coresets found by our approach to downstream models with different architectures including those based on convolutional neural networks, ResNets and ViTs. Results in Table 3 and Table 13 (in App. C.4) show that our coresets achieve superior performance than random for all the architectures. Thus, our concept-based score in conjunction with stratified sampling zheng2022coverage which balances the number of easy/hard examples in the coreset is an effective approach for both standard and label-free CS at various pruning rates.
@@
5.2 Evaluating efficiency of CS via our concept-based approach
Here we compare the efficiency of our approach at finding coresets compared to approaches based on training dynamics. For training dynamics-based approaches the time required to find the coreset is dominated by the time needed to train the downstream model on the entire dataset first. For example, finding a coreset of Imagenet using a ResNet-34 model takes roughly 8 hours on two A-100 GPUs. In comparison, for our approach, extracting concepts via LLMs (block 1 of Fig. 2) takes 3 seconds per class (totaling to 25 minutes for all classes of Imagenet using 2 GPUs), and computing visual/concept features via CLIP and training a linear concept bottleneck layer (block 2 for Fig. 2) takes roughly 30 minutes for Imagenet offering $\approx 8$x speedup for CS over approaches relying on training dynamics of the downstream model. Moreover, since our method is independent of the downstream model architecture we do not need to repeat the CS step for different architectures. This is in contrast with training dynamics-based methods that require training the downstream model for every new architecture to find the corresponding best coreset for it.
Table 3: Superior performance of downstream models with various architectures trained on our coresets for Imagenet compared to Random for standard (Ours) and label-free (Ours-LF) CS highlight our method’s effectiveness for CS without knowledge of the downstream model.
Model Architecture Method Pruning Rates
@@ 5.2 Evaluating efficiency of CS via our concept-based approach
RN-18 Random $71.15_{\scriptscriptstyle\pm 0.23}$ $68.48_{\scriptscriptstyle\pm 0.10}$ $63.15_{\scriptscriptstyle\pm 0.19}$ $44.96_{\scriptscriptstyle\pm 0.50}$
Ours-LF $71.21_{\scriptscriptstyle\pm 0.09}$ $68.77_{\scriptscriptstyle\pm 0.13}$ $63.76_{\scriptscriptstyle\pm 0.09}$ $47.50_{\scriptscriptstyle\pm 0.10}$
Ours $70.94_{\scriptscriptstyle\pm 0.19}$ $69.30_{\scriptscriptstyle\pm 0.84}$ $65.16_{\scriptscriptstyle\pm 0.04}$ $49.57_{\scriptscriptstyle\pm 0.16}$
RN-50 Random $76.06_{\scriptscriptstyle\pm 0.11}$ $74.44_{\scriptscriptstyle\pm 0.04}$ $70.50_{\scriptscriptstyle\pm 0.02}$ $53.56_{\scriptscriptstyle\pm 0.13}$
Ours-LF $76.54_{\scriptscriptstyle\pm 0.08}$ $74.84_{\scriptscriptstyle\pm 0.03}$ $71.10_{\scriptscriptstyle\pm 0.09}$ $55.22_{\scriptscriptstyle\pm 0.92}$
Ours $76.29_{\scriptscriptstyle\pm 0.10}$ $75.12_{\scriptscriptstyle\pm 0.03}$ $72.05_{\scriptscriptstyle\pm 0.09}$ $58.26_{\scriptscriptstyle\pm 0.46}$
DenseNet Random $76.55_{\scriptscriptstyle\pm 0.11}$ $75.31_{\scriptscriptstyle\pm 0.15}$ $71.85_{\scriptscriptstyle\pm 0.20}$ $56.27_{\scriptscriptstyle\pm 0.21}$
Ours-LF $76.94_{\scriptscriptstyle\pm 0.23}$ $75.52_{\scriptscriptstyle\pm 0.11}$ $72.28_{\scriptscriptstyle\pm 0.14}$ $58.16_{\scriptscriptstyle\pm 0.04}$
Ours $76.80_{\scriptscriptstyle\pm 0.22}$ $75.94_{\scriptscriptstyle\pm 0.03}$ $73.40_{\scriptscriptstyle\pm 0.27}$ $60.93_{\scriptscriptstyle\pm 0.19}$
Wide Resnet Random $77.39_{\scriptscriptstyle\pm 0.19}$ $75.32_{\scriptscriptstyle\pm 0.12}$ $71.51_{\scriptscriptstyle\pm 0.10}$ $54.78_{\scriptscriptstyle\pm 0.83}$
Ours-LF $77.83_{\scriptscriptstyle\pm 0.08}$ $75.77_{\scriptscriptstyle\pm 0.09}$ $71.97_{\scriptscriptstyle\pm 0.08}$ $57.34_{\scriptscriptstyle\pm 0.58}$
Ours $77.48_{\scriptscriptstyle\pm 0.05}$ $76.32_{\scriptscriptstyle\pm 0.03}$ $73.14_{\scriptscriptstyle\pm 0.03}$ $59.75_{\scriptscriptstyle\pm 0.70}$
5.3 Visualizing easy and challenging samples based on our concept-based score